# DISCOVER: Automated Curricula for Sparse-Reward Reinforcement Learning

### v1

@@ Acknowledgments
[…]
0[2]
@@ Appendix A Additional Related Work
DISCOVER is designed to efficiently address the exploration-exploitation dilemma, a key concept in sequential decision making. Fundamentally, it expresses an agent’s inevitable trade-off between the objectives of learning its environment and solving a task. A particularly common approach to balancing exploration and exploitation is grounded in the principle of optimism in the face of uncertainty [e.g., 69]. Thereby, the agent selects actions that maximize an upper confidence bound (UCB) of the reward function, i.e., it selects actions which based on the agents’ imperfect knowledge could lead to a large reward. In many settings of sequential decision making, such as linear bandits, this approach achieves the rate-optimal regret $\smash{R_{T}\sim\widetilde{O}(d\sqrt{T})}$ [69, 2, 1, 13]. The DISCOVER objective extends UCB to the problem of goal selection in RL. Beyond UCB, many other methods have been shown to effectively direct exploration towards “relevant” experience, such as in bandits [e.g., 60, 59, 34], in RL [e.g., 14, 73], or in active learning [e.g., 42, 37, 6].
@@ Appendix B Connection to the Likelihood of Success
for arbitrary $s_{0}\in\mathcal{S},g,g^{\star}\in\mathcal{G}$. Note that the direction of the inequality is flipped as the value function represents a negative distance. For $\alpha=0.5$, DISCOVER maximizes a probabilistic estimate of the left hand side of Equation 5, which is a tight lower-bound to $V^{\pi}(s_{0},g^{\star})$. The value $V^{\pi}(s_{0},g^{\star})$ is exactly the quantity of interest for the agent, as it represents the negative expected number of steps to reach the true goal $g^{\star}$. Intuitively, DISCOVER selects goals that are optimistically going to guarantee the shortest path to the actual goal or, in the undiscounted case, the likelihood of reaching it within an episode.
@@ C.2 Assumptions
$V_{t}(s_{0},\cdot)$ =$\overset{\mathrm{def}}{=}\phi(\cdot)^{\top}(\Sigma_{t}+\lambda I)^{-1}\sum_{s=h}^{t-1}\phi(g_{s})y_{s},$
$\sigma_{t}(s_{0},\cdot)$ =$\overset{\mathrm{def}}{=}\sqrt{\phi(\cdot)^{\top}(\Sigma_{t}+\lambda I)^{-1}\phi(\cdot)},$
@@
Relaxing Assumption C.2
One can consider any individual feedback $y_{t}$ as being the result of an oracle that achieves the commanded goal $g_{t}$ with probability at least $\alpha$ within $K$ episodes. With this looser assumption, our bound in Informal Theorem 4.2 simply increases by a factor $K$.
C.3 Proof of Informal Theorem 4.2
@@
Let Assumptions C.1, C.2 and C.3 hold. Fix any $\delta\in(0,1),n\geq 1,\alpha\in(0,1)$, and let
@@
By Theorem 3 in [13] and using that the feature spaces $\phi$ and $\varphi$ are orthogonal (cf. Assumption C.1), we have
@@
with $B=1$. Bounding $\gamma_{T}\leq O(d\log T)$ using Assumption C.1, completes the proof. ∎
@@
Let Assumption C.4 hold and fix any $\epsilon>0,\alpha>0$. Then, for all $g^{\prime}\in\mathcal{G}$ with $d^{\star}(g^{\prime},g^{\star})\geq\epsilon$ there exists a $g\in\mathcal{G}$ with $d^{\star}(g,g^{\prime})=\epsilon$ such that $r^{\star}(g)-r^{\star}(g^{\prime})\geq(1-2\alpha)\epsilon$.
@@
By Assumption C.4, for any $d^{\star}(g^{\prime},g)\in[0,\epsilon]$, we have that $g\in\mathcal{G}$. We take the goal $g$ such that $d^{\star}(g^{\prime},g)=\epsilon$. We then obtain
@@
Let Assumptions C.1, C.2, C.3 and C.4 hold with $\beta_{t}$ as in Equation 6. Fix any ${\delta\in(0,1)}$, ${n\geq 1}$, ${\epsilon>0}$, ${\alpha\in(0,\frac{1}{2})}$ and $0<\Delta<(1-2\alpha)\epsilon$. With probability $1-\delta$, there exist a ${t^{\prime}\in\{n,\dots,n+T\}}$ with $\smash{T=\widetilde{\Theta}(\frac{d^{2}}{\alpha((1-2\alpha)\epsilon-\Delta)^{2}})}$ and a $\tilde{g}\in\mathcal{G}$ with $d^{\star}(g_{t^{\prime}},\tilde{g})\leq\epsilon$ such that
@@
With probability $1-\frac{\delta}{2}$, the number of episodes $T$ until the agent has achieved $T_{\mathrm{ach}}$ of its goals is bounded as ${T=\widetilde{\Theta}(\frac{T_{\mathrm{ach}}}{\alpha})}$. [^5] See Lemma C.10. We denote by $\mathcal{T}\subseteq\{n,\dots,n+T\}$ the set of episodes in which the agent has achieved its goal. Further, by Proposition C.6, also with probability $1-\frac{\delta}{2}$, we have
@@
Further, by Lemma C.7, there exists a $\tilde{g}\in\mathcal{G}$ such that $d^{\star}(g_{t^{\prime}},\tilde{g})=\epsilon$ and
@@
Figure 8: Plotting the effect of the parameter $\alpha$ on the (inverse) goal-achieval rate (cf. Theorem C.9). The target goal is reached fastest for $\alpha\approx 0.1$.
@@
Let Assumptions C.1, C.2, C.3, C.4 and C.5 hold with $\beta_{t}$ as in Equation 6. Fix any ${\delta\in(0,1)}$, ${\alpha\in(0,\frac{1}{2})}$, and let $D=d^{\star}(s_{0},g^{\star})$. Then, with probability $1-\delta$, selecting goals $g_{t}$ with DISCOVER($\alpha,\beta_{t}$), the number of episodes $N$ until $g^{\star}\in\mathcal{G}_{N}$ is bounded by $N\leq\widetilde{O}(\frac{Dd^{2}}{\alpha(1-2\alpha)^{2}(1-\alpha)^{3}\kappa^{3}})=\widetilde{O}(\frac{Dd^{2}}{\kappa^{3}})$.
@@
We prove the theorem by applying Lemma C.8 $M\overset{\mathrm{def}}{=}\lceil\frac{(1-2\alpha)D}{\Delta}\rceil$ times, while setting ${\epsilon=(1-\alpha)\kappa}$. First, for an arbitrary $10\leq i\leq M-$ 1, we assume for the goal set $\mathcal{G}_{iT}$ with some $\smash{T=\widetilde{\Theta}(\frac{d^{2}}{\alpha((1-2\alpha)(1-\alpha)\kappa-\Delta)^{2}})}$ that it holds that
@@
Now, applying Lemma C.8 yields that after an additional $T$ steps, with high probability, there exists a $t^{\prime}\in\{iT,\dots,(i+1)T\}$ such that there is a $\tilde{g}\in\mathcal{G}$ with $d^{\star}(g_{t^{\prime}},\tilde{g})\leq(1-\alpha)\kappa$ satisfying
@@
Hence, by Assumption C.5, we have that $\tilde{g}\in\mathcal{G}_{(i+1)T}$, and therefore,
@@
Finally, we include a technical lemma that is used in the proof of Lemma C.8.
@@
Let $0<\alpha\leq 1$, $10<\delta<$ 1 and suppose $T_{\mathrm{ach}}\geq 8\log(1/\delta)$. Furthermore, let ${X_{1},\dots,X_{T}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Bern}(\alpha)}$ and $S_{T}=\sum_{t=1}^{T}X_{t}$. Then, for some $T=\widetilde{\Theta}(\tfrac{T_{\mathrm{ach}}}{\alpha})$, with probability $1-\delta$, we have $S_{T}\geq T_{\mathrm{ach}}$.
@@
Note that $\mu\geq(1+\gamma)^{2}T_{\mathrm{ach}}$ and $10<\gamma<$ 1 (for $T_{\mathrm{ach}}\geq 8\log(1/\delta)$). Hence,
@@ Appendix D Additional Experimental Results
In this section, we present additional experiments and ablations. The environments used for this evaluation are visualized in Figure 9 [9].
@@
In complex environments, obtaining accurate direction estimates can be challenging. To evaluate the utility of directed goal selection in scenarios where direction estimates are imprecise, we add Gaussian noise to the target goal location. We then compare the number of environment steps required to reach a $10\%$ target goal achievement rate using the hand-designed goal selection strategy from Figure 6, under varying levels of noise variance. The results in Figure 10 show that even with substantial noise, pointmaze environments that are unsolvable by undirected methods, remain solvable by DISCOVER. This demonstrates that even imprecise directional estimates can significantly aid target goal discovery in complex goal spaces.
@@
In Figure 11, we evaluate the effect of using the simple proposed adaptation strategy with different target goal achievement ratio $p_{A}^{\star}$, and compare with fixing the $\alpha_{t}$ parameters to 00 (Target Relevance + Novelty) and $0.5$ (Fixed DISCOVER). Furthermore, we report the average $\alpha_{t}$ for DISCOVER for all easy and hard tasks respectively in Figure 12. The comparison of the success rates on the antmaze environments demonstrates that the adaptation strategy with any target goal achievement works better than fixing the parameters. This can be observed from the goal achievement rates. If we fix $\alpha_{t}=0.5$, we choose goals that are "too easy" and therefore don’t explore sufficiently. On the other hand, by fixing $\alpha_{t}=0$ we select goals that are "too hard", which also leads too poor improvement. By using the simple adaptation strategy, we roughly achieve the target goal achievement specified. The optimal performance is achieved for the target goal achievement $p_{A}^{\star}=0.5$, which is in line with what other methods found [53, 41, 85]. The average $\alpha_{t}$, which is found by the adaptation strategy, initially goes up to $0.15$, which is roughly what we found in the theoretical analysis in the linear bandit setting (cf. Theorem C.9), and then starts to decay. The decay can be explained by the fact that once we can reach the target goal, we don’t need to optimize for achievability anymore.
@@
In Figure 13, we study how the standard deviation $\sigma(g,g^{\star})$ from goal $g$ to the target $g^{\star}$ influences the training. This term theoretically is part of the UCB term, directing the agent towards the target goal. To this end, we fix the contribution of $\sigma(s_{0},g)$ (i.e., set $\beta_{t}=\frac{1}{\alpha_{t}}$) and consider a separate fixed $\beta_{t}^{\prime}=\frac{\beta^{\prime}}{1-\alpha}$, which determines the contribution of $\sigma(g,g^{\star})$. The plot shows that no value for the $\beta^{\prime}$ parameter has a significant positive effect on the performance. For this reason, we substitute $\sigma(s_{0},g)$ for $\sigma(g,g^{\star})$ in our other experiments with DISCOVER.
@@
We visualize the different components of the DISCOVER objective over the course of training in Figure 15. The first term $V(s_{0},g)$ has high-value close to the initial state. By maximizing it, we will pick a goal that is close to the start and likely to be achievable, which matches the intuition. The second term $V(g,g^{\star})$ represents the value from a goal $g$ to the target goal $g^{\star}$. The plots show that in the first episodes the value is small everywhere and only once goals are discovered that are closer to the final target we observe higher values. Over the course of training, its role of encouraging to pick goals relevant to the final target becomes more evident. This term therefore directs the goal selection towards the final goal. Finally, the standard deviation $\sigma(s_{0},g)$ has the largest value at the border of the current achievable goal set and therefore encourages selecting novel goals. In general, the components of the DISCOVER objective during the training match the previously presented intuition and can efficiently guide the goal selection towards the desired target.
@@
Figure 16: Visualization of the exploration of the standard non-goal-conditioned RL methods in the antmaze environments. We run MaxInfoRL [72] with SAC.

### v2

@@ Acknowledgments
appendix.Aappendix.Aappendix.Aappendix.Aappendix.Bappendix.Csubsection.C.1subsection.C.2theorem.C.5subsection.C.3appendix.Dfigure.caption.17figure.caption.17figure.caption.19figure.caption.21figure.caption.21figure.caption.24figure.caption.24appendix.Esubsection.E.1subsection.E.2subsection.E.3
@@ Appendix A Additional Related Work
DISCOVER is designed to efficiently address the exploration-exploitation dilemma, a key concept in sequential decision making. Fundamentally, it expresses an agent’s inevitable trade-off between the objectives of learning its environment and solving a task. A particularly common approach to balancing exploration and exploitation is grounded in the principle of optimism in the face of uncertainty [e.g., 68]. Thereby, the agent selects actions that maximize an upper confidence bound (UCB) of the reward function, i.e., it selects actions which based on the agents’ imperfect knowledge could lead to a large reward. In many settings of sequential decision making, such as linear bandits, this approach achieves the rate-optimal regret $\smash{R_{T}\sim\widetilde{O}(d\sqrt{T})}$ [68, 2, 1, 13]. The DISCOVER objective extends UCB to the problem of goal selection in RL. Beyond UCB, many other methods have been shown to effectively direct exploration towards “relevant” experience, such as in bandits [e.g., 60, 59, 34], in RL [e.g., 14, 72], or in active learning [e.g., 42, 37, 6].
@@ Appendix B Connection to the Likelihood of Success
for arbitrary $s_{0}\in\mathcal{S},g,g^{\star}\in\mathcal{G}$. Note that the direction of the inequality is flipped as the value function represents a negative distance. For $\alpha=0.5$, DISCOVER maximizes a probabilistic estimate of the left hand side of Equation˜5, which is a tight lower-bound to $V^{\pi}(s_{0},g^{\star})$. The value $V^{\pi}(s_{0},g^{\star})$ is exactly the quantity of interest for the agent, as it represents the negative expected number of steps to reach the true goal $g^{\star}$. Intuitively, DISCOVER selects goals that are optimistically going to guarantee the shortest path to the actual goal or, in the undiscounted case, the likelihood of reaching it within an episode.
@@ C.2 Assumptions
$V_{t}(s_{0},\cdot)$ $\overset{\mathrm{def}}{=}\phi(\cdot)^{\top}(\Sigma_{t}+\lambda I)^{-1}\sum_{s=h}^{t-1}\phi(g_{s})y_{s},$
$\sigma_{t}(s_{0},\cdot)$ $\overset{\mathrm{def}}{=}\sqrt{\phi(\cdot)^{\top}(\Sigma_{t}+\lambda I)^{-1}\phi(\cdot)},$
@@
Relaxing Assumption˜C.2
One can consider any individual feedback $y_{t}$ as being the result of an oracle that achieves the commanded goal $g_{t}$ with probability at least $\alpha$ within $K$ episodes. With this looser assumption, our bound in Informal Theorem˜4.2 simply increases by a factor $K$.
C.3 Proof of Informal Theorem˜4.2
@@
Let Assumptions˜C.1, C.2 and C.3 hold. Fix any $\delta\in(0,1),n\geq 1,\alpha\in(0,1)$, and let
@@
By Theorem 3 in [13] and using that the feature spaces $\phi$ and $\varphi$ are orthogonal (cf. Assumption˜C.1), we have
@@
with $B=1$. Bounding $\gamma_{T}\leq O(d\log T)$ using Assumption˜C.1, completes the proof. ∎
@@
Let Assumption˜C.4 hold and fix any $\epsilon>0,\alpha>0$. Then, for all $g^{\prime}\in\mathcal{G}$ with $d^{\star}(g^{\prime},g^{\star})\geq\epsilon$ there exists a $g\in\mathcal{G}$ with $d^{\star}(g,g^{\prime})=\epsilon$ such that $r^{\star}(g)-r^{\star}(g^{\prime})\geq(1-2\alpha)\epsilon$.
@@
By Assumption˜C.4, for any $d^{\star}(g^{\prime},g)\in[0,\epsilon]$, we have that $g\in\mathcal{G}$. We take the goal $g$ such that $d^{\star}(g^{\prime},g)=\epsilon$. We then obtain
@@
Let Assumptions˜C.1, C.2, C.3 and C.4 hold with $\beta_{t}$ as in Equation˜6. Fix any $\delta\in(0,1)$}, $n\geq 1$}, $\epsilon>0$}, $\alpha\in(0,\frac{1}{2})$} and $0<\Delta<(1-2\alpha)\epsilon$. With probability $1-\delta$, there exist a $t^{\prime}\in\{n,\dots,n+T\}$} with $\smash{T=\widetilde{\Theta}(\frac{d^{2}}{\alpha((1-2\alpha)\epsilon-\Delta)^{2}})}$ and a $\tilde{g}\in\mathcal{G}$ with $d^{\star}(g_{t^{\prime}},\tilde{g})\leq\epsilon$ such that
@@
With probability $1-\frac{\delta}{2}$, the number of episodes $T$ until the agent has achieved $T_{\mathrm{ach}}$ of its goals is bounded as $T=\widetilde{\Theta}(\frac{T_{\mathrm{ach}}}{\alpha})$}. [^5] See Lemma C.10. We denote by $\mathcal{T}\subseteq\{n,\dots,n+T\}$ the set of episodes in which the agent has achieved its goal. Further, by Proposition˜C.6, also with probability $1-\frac{\delta}{2}$, we have
@@
Further, by Lemma˜C.7, there exists a $\tilde{g}\in\mathcal{G}$ such that $d^{\star}(g_{t^{\prime}},\tilde{g})=\epsilon$ and
@@
Figure 7: Plotting the effect of the parameter $\alpha$ on the (inverse) goal-achieval rate (cf. Theorem˜C.9). The target goal is reached fastest for $\alpha\approx 0.1$.
@@
Let Assumptions˜C.1, C.2, C.3, C.4 and C.5 hold with $\beta_{t}$ as in Equation˜6. Fix any $\delta\in(0,1)$}, $\alpha\in(0,\frac{1}{2})$}, and let $D=d^{\star}(s_{0},g^{\star})$. Then, with probability $1-\delta$, selecting goals $g_{t}$ with DISCOVER($\alpha,\beta_{t}$), the number of episodes $N$ until $g^{\star}\in\mathcal{G}_{N}$ is bounded by $N\leq\widetilde{O}(\frac{Dd^{2}}{\alpha(1-2\alpha)^{2}(1-\alpha)^{3}\kappa^{3}})=\widetilde{O}(\frac{Dd^{2}}{\kappa^{3}})$.
@@
We prove the theorem by applying Lemma˜C.8 $M\overset{\mathrm{def}}{=}\lceil\frac{(1-2\alpha)D}{\Delta}\rceil$ times, while setting $\epsilon=(1-\alpha)\kappa$}. First, for an arbitrary $0\leq i\leq M-1$, we assume for the goal set $\mathcal{G}_{iT}$ with some $\smash{T=\widetilde{\Theta}(\frac{d^{2}}{\alpha((1-2\alpha)(1-\alpha)\kappa-\Delta)^{2}})}$ that it holds that
@@
Now, applying Lemma˜C.8 yields that after an additional $T$ steps, with high probability, there exists a $t^{\prime}\in\{iT,\dots,(i+1)T\}$ such that there is a $\tilde{g}\in\mathcal{G}$ with $d^{\star}(g_{t^{\prime}},\tilde{g})\leq(1-\alpha)\kappa$ satisfying
@@
Hence, by Assumption˜C.5, we have that $\tilde{g}\in\mathcal{G}_{(i+1)T}$, and therefore,
@@
Finally, we include a technical lemma that is used in the proof of Lemma˜C.8.
@@
Let $0<\alpha\leq 1$, $0<\delta<1$ and suppose $T_{\mathrm{ach}}\geq 8\log(1/\delta)$. Furthermore, let $X_{1},\dots,X_{T}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Bern}(\alpha)$} and $S_{T}=\sum_{t=1}^{T}X_{t}$. Then, for some $T=\widetilde{\Theta}(\tfrac{T_{\mathrm{ach}}}{\alpha})$, with probability $1-\delta$, we have $S_{T}\geq T_{\mathrm{ach}}$.
@@
Note that $\mu\geq(1+\gamma)^{2}T_{\mathrm{ach}}$ and $0<\gamma<1$ (for $T_{\mathrm{ach}}\geq 8\log(1/\delta)$). Hence,
@@ Appendix D Additional Experimental Results
In this section, we present additional experiments and ablations. The environments used for this evaluation are visualized in Figure˜8 [9].
@@
In complex environments, obtaining accurate direction estimates can be challenging. To evaluate the utility of directed goal selection in scenarios where direction estimates are imprecise, we add Gaussian noise to the target goal location. We then compare the number of environment steps required to reach a $10\%$ target goal achievement rate using the hand-designed goal selection strategy from Figure˜5, under varying levels of noise variance. The results in Figure˜9 show that even with substantial noise, pointmaze environments that are unsolvable by undirected methods, remain solvable by DISCOVER. This demonstrates that even imprecise directional estimates can significantly aid target goal discovery in complex goal spaces.
@@
In Figure˜10, we evaluate the effect of using the simple proposed adaptation strategy with different target goal achievement ratio $p_{A}^{\star}$, and compare with fixing the $\alpha_{t}$ parameters to 0 (Target Relevance + Novelty) and $0.5$ (Fixed DISCOVER). Furthermore, we report the average $\alpha_{t}$ for DISCOVER for all easy and hard tasks respectively in Figure 11. The comparison of the success rates on the antmaze environments demonstrates that the adaptation strategy with any target goal achievement works better than fixing the parameters. This can be observed from the goal achievement rates. If we fix $\alpha_{t}=0.5$, we choose goals that are "too easy" and therefore don’t explore sufficiently. On the other hand, by fixing $\alpha_{t}=0$ we select goals that are "too hard", which also leads too poor improvement. By using the simple adaptation strategy, we roughly achieve the target goal achievement specified. The optimal performance is achieved for the target goal achievement $p_{A}^{\star}=0.5$, which is in line with what other methods found [53, 41, 84]. The average $\alpha_{t}$, which is found by the adaptation strategy, initially goes up to $0.15$, which is roughly what we found in the theoretical analysis in the linear bandit setting (cf. Theorem˜C.9), and then starts to decay. The decay can be explained by the fact that once we can reach the target goal, we don’t need to optimize for achievability anymore.
@@
In Figure˜12, we study how the standard deviation $\sigma(g,g^{\star})$ from goal $g$ to the target $g^{\star}$ influences the training. This term theoretically is part of the UCB term, directing the agent towards the target goal. To this end, we fix the contribution of $\sigma(s_{0},g)$ (i.e., set $\beta_{t}=\frac{1}{\alpha_{t}}$) and consider a separate fixed $\beta_{t}^{\prime}=\frac{\beta^{\prime}}{1-\alpha}$, which determines the contribution of $\sigma(g,g^{\star})$. The plot shows that no value for the $\beta^{\prime}$ parameter has a significant positive effect on the performance. For this reason, we substitute $\sigma(s_{0},g)$ for $\sigma(g,g^{\star})$ in our other experiments with DISCOVER.
@@
We visualize the different components of the DISCOVER objective over the course of training in Figure˜14. The first term $V(s_{0},g)$ has high-value close to the initial state. By maximizing it, we will pick a goal that is close to the start and likely to be achievable, which matches the intuition. The second term $V(g,g^{\star})$ represents the value from a goal $g$ to the target goal $g^{\star}$. The plots show that in the first episodes the value is small everywhere and only once goals are discovered that are closer to the final target we observe higher values. Over the course of training, its role of encouraging to pick goals relevant to the final target becomes more evident. This term therefore directs the goal selection towards the final goal. Finally, the standard deviation $\sigma(s_{0},g)$ has the largest value at the border of the current achievable goal set and therefore encourages selecting novel goals. In general, the components of the DISCOVER objective during the training match the previously presented intuition and can efficiently guide the goal selection towards the desired target.
@@
Figure 15: Visualization of the exploration of the standard non-goal-conditioned RL methods in the antmaze environments. We run the state-of-the-art MaxInfoRL [71] with SAC.
(a) Direction estimates become useful before the final goal is reached. We plot the pearson correlation between the current relevance estimate and the L2-distance between the goal and the target goal for the hard variants of the antmaze and pointmaze environments.
(b) A value function ensemble size of six strikes a good balance between computational cost and the quality of uncertainty estimates, leading to improved exploration and task performance. We plot the performance of DISCOVER varying the number of value function ensemble members in the hard antmaze environment.
Investigation of Direction Estimation
To further investigate the ability of the relevance term in the DISCOVER objective to direct the exploration towards to the target goal, we plot the pearson correlation between the relevance term $V^{\star}(g,g^{\star})$ and a proxy for the true distance $g-g^{\star}||$ in Figure˜16(a). The correlation in both the antmaze and pointmaze environments increases quickly during training. This indicates that the relevance term captures a useful notion of distance, even before the final goal is reached for the first time. This behavior explains DISCOVER’s superior performance compared to undirected goal-selection strategies, as the value function is able to capture a useful notion of distance to the target, before it was reached. This allows DISCOVER to effectively guide the exploration to the true target goal. Notably, the correlation does not converge to one. This is expected, as the Euclidean distance in the goal space is an imperfect proxy for the true shortest traversable path within the environment.
Ablation of Ensemble Size
A key parameter in the DISCOVER algorithm is the value function ensemble size. To quantify the impact of ensemble size on uncertainty estimates, we evalute the performance of DISCOVER in the hard antmaze environment, comparing ensembles of 2, 4, 6, and 8 critics in Figure˜16(b). Consistent with prior work [48], increasing the number of critics above the default of two improves uncertainty estimation. In our experiments, an ensemble of six critics proved sufficient to yield robust performance.
Relying on an ensemble of critics can be computationally expensive. We find that in our experiments the overhead of increasing the critic size from 2 to 6 was modest. This is mainly due to small critic networks and highly parallel training using jax. In practice, we find that increasing the ensemble from 2 to 6 critics was sufficient to yield high‐quality uncertainty estimates, while incurring only a  26% increase in runtime.